Please help me this problem:
Write an equation for the hyperbola with vertices at (-3,0) and (3,0) and passing through (6,1)
Thanks

Answer by stanbon(75887)

Please help me this problem:
Write an equation for the hyperbola with vertices at (-3,0) and (3,0) and passing through (6,1)
-------
Since the vertices are symmetric to (0,0), the center is at the origin.
-----
Form:
[x^2/a^2] - [y^2/b^2] = 1
---
Substitute and solve for "b^2":
[6^2/3^2 - [1^2/b]^2 = 1
----
4 - [1/b^2] = 1
---
(1/b^2) = 3
b^2 = 1/3
------
Equation:
[x^2/9] - [y^2/(1/3)] = 1
-----
Cheers,
Stan H.